### v3

@@ 4 Mechanics of free growth
5 Why we test in acceleration alone
We agree with West, Ricardo and Malthus, and most economists since, that innovation is a prerequisite of growth. We go further, and expect that substantially all capital growth is explained by the innovator’s reserve, so that net saving is obviated. This prediction, however, cannot be tested ideally from equations for growth as distinct from acceleration. In Eq. (5), for example, where thrift theory predicts $E(q(K)_{i})=0$, free growth theory does not predict the diametric opposite $E(s(K)_{i})=0$. Free growth theory makes no prediction as to the amount of saving or investment, or of its ratio to capital or to capital growth, but rather questions its effect on capital growth. Causality is more clearly revealed in capital acceleration, where balanced growth is impossible, and where changes in thrift can thus be compared to concurrent changes in growth. The only predictions of free growth theory which can be tested to give unambiguous results are Eqs. (11) and (19), which predict in capital acceleration alone.
@@ 8 Data sources
All our data are drawn from Distributional National Accounts (DINA) from the free online database World Inequality Database (WID). This source collates data from national accounts and tax data of 105 countries in constant currency units, and adjusts them where needed to conform to current standards of the System of National Accounts (SNA) published by the United Nations. We show results for the 88 of those countries which report all three of the factors, namely net saving, consumption and market-value capital, needed for deriving the thrift and free growth indexes. WID’s source for these data is national accounts.
Consumption $C$ in our text and equations is reproduced from Final Consumption Expenditure (mcongo) [^5] WID code. This sums personal consumption expenditure PCE and government expenditure GCE. Net saving $S_{net}$ and market-value $K$ are taken from net national saving (msavin) and market-value Capital Wealth (mnweal) respectively. GDP, which we use only for weighting purposes in Figs. 1 and 2, is reproduced from GDP (mgdpro).
@@ 9 Accessing our results and methods
Tables and other displays of our findings for each country, and showing our methods of calculation, can be accessed at the web appendix (https://3woilz-0-0.shinyapps.io/RhinoApplication/).
@@ 10 Displays
Eqs. (10) and (18) show $\theta_{s}+\varphi_{s}=1$ and $\theta_{c}+\varphi_{c}=1$. All displays here and in the web appendix, except for Figs. 1 and 2 and Tables 1 and 3, save space by showing $\varphi_{s}$ and $\varphi_{c}$ only, leaving $\theta_{s}$ and $\theta_{c}$ implicit as their complements to unity. Table 2 shows $\varphi_{s}$ and $\varphi_{c}$ for each of the 88 countries averaged over all years.
The web appendix includes displays of the $\varphi_{s}$ and $\varphi_{c}$ for each year in each country over the report period. These tend to show upward and downward spikes in values of $\varphi_{s}$ and $\varphi_{c}$ in some years. Those spikes tend to be associated with small absolute values of denominators, in these cases $\Delta g(K)$, in those countries and years. Small denominators magnify errors in measurements of numerators. Worse, when $\Delta g(K)_{i}$ is small, small mismeasurements of $g(K)_{i}$ or $g(K)_{i-1}$ might reverse $\Delta g(K)_{i}$ in sign.
To maximize reliability of test results, we apply a range of screens to omit years where absolute denominators fall below a given threshold. Some displays show $\varphi_{s}$ and $\varphi_{c}$ for all years, regardless of denominator size. Others screen out all years where absolute denominators are less then .01, then .025, then continuing upward in increments of .025 to a maximum screen of .15. $\varphi_{s}$ and $\varphi_{c}$ are plotted for each country unscreened and at each of the seven successive levels of screening. Figs. 1 and 2, and all four Tables, applied a screen of .01. The denominator whose absolute value is screened is capital acceleration $\Delta g(K)$ in all displays.
@@
Readers are encouraged to experiment with the interactive web appendix, for example in changing screen levels to see effects on $\varphi$ and $\theta$.
@@ 12 Discussion and conclusions
These findings challenge the teachings that capital growth is effected by net saving enabled by consumption restraint, and that producer cost, including imputed interest as the opportunity cost of capital, converges to market realization. Evidence showing $\overline{\varphi_{s,i}}\cong 1$ and $\overline{\varphi_{c,i}}\cong 1$ suggests that all capital growth is free, and consequently that market realization, in the presence of innovation, exceeds producer cost by the entirety of capital growth. Meanwhile the same evidence, which indicates that net saving adds no capital value, suggests a review of the teachings that consumption plus net saving gives net income, and that consumption plus net investment gives net output.
@@ Appendix B.  Net output with human capital
It is possible in principle to model a value-added chain which includes human capital, and to compare findings with those shown in Eq. (B.1). Let human capital $H$, in that new model, stand as the last link in the value-added chain. Adapting the classic illustration of the value added principle, say that farms produce wheat, mills convert the wheat to flour, bakeries convert the flour into bread, and humans convert some of the bread, called invested consumption, into human capital. The net outflow from this extended value-added chain is not all of consumption, but only pure consumption, meaning the part remaining after the part invested in human capital (invested consumption) is subtracted [^7] The concepts of invested consumption, self-invested work and human depreciation were introduced in Schultz (1961).. By this reasoning, the principle that net output is expressed in capital growth plus net outflow gives
@@
if Schultz and Ben-Porath are right.
@@ Appendix C.  Mill’s statement of the free growth idea
This passage may be the first clear statement of what we call free growth theory. His use of the words "accelerate" and "concurrently" suggest that his path of reasoning was something like ours, and that he meant what we call free growth rather than the alternating phases of higher and lower saving rates pictured in thrift theory. His use of the word "saving" is logical, but is not the usual one followed by us and used in national accounts.

### v4

@@ 4 Mechanics of free growth
5 Testing directly from capital growth
The essential tenet of thrift theory shows in Eq. (5), which expects capital growth to equal net investment within the authorized rate. We have first tested it, and its counterpart in Eq. (13), from their implications for capital acceleration, rather than directly, because causalities connecting variables tend to be most clearly revealed in simultaneous changes in those variables. In this case, causalities connected to $g(K)$ are likeliest to be revealed in changes to $g(K)$, meaning capital acceleration $\Delta g(K)$.
Results of those tests were vivid. Tables and figures above seem to show that change in net investment, expected in thrift theory to explain all of capital acceleration, explains none of it. We will now test Eqs. (5) and (13) directly, but we cannot expect equal clarity in test results. Free growth theory predicts that capital growth is explained by the innovator’s reserve alone, and is not a function of $s^{*}$ or $c^{*}$. For this reason, it makes no predictions as to the relationships among $g(K),s^{*}$ and $c^{*}$ at any time. It expects rather that Eq. (13) might predict correctly in occasional coincidence, but not in general. By Eq. (12), versions of the thrift index applying to capital growth directly may be given as
$\theta^{*}_{s,i}=\frac{s^{*}_{i}}{g(K)_{i}}\ ,\quad\theta^{*}_{c,i}=\frac{(1-c^{*})_{i}}{g(K)_{i}}\ ,\quad\theta^{*}_{s}=E(\theta^{*}_{s,i})\quad\text{and}\quad\theta^{*}_{c}=E(\theta^{*}_{c,i})$
Eq. (13) now implies
$E(\theta^{*}_{s})=E(\theta^{*}_{c})=1\quad\text{under thrift assumptions.}$. (21)
Again, free growth theory makes no prediction as to the relationship of capital growth to net saving or to consumption.
Fig. 3 plots average $\theta^{*}_{s}$ and $\theta^{*}_{c}$, GDP-weighted, for 86 countries over the period 1980-2022. Table 4 shows associated regressions. Table 5 shows $\theta^{*}_{s}$ and $\theta^{*}_{c}$ for each of these 86 countries for which data were reported. As with tests from capital acceleration, these findings do not seem to support thrift theory.
@@ 6 Optimum investment policy
Fig. 3: $\theta^{*}_{s}$ and $\theta^{*}_{c}$ (86 countries)
Table 4: Regression of $s^{*}$ and $1-c^{*}$ on $g(K)$ (Screen = 0.01). $H_{0}$ per thrift theory: $g(K)\cong s^{*}$ and $g(K)\cong 1-c^{*}$.
$s^{*}$ $1-c^{*}$
Regression of value shown on $g(K)$ 0.2103∗∗∗ -0.6760∗∗∗
(0.0541) (0.1480)
Observations 1,691 1,665
R2 0.76527 0.83560
Within R2 0.14579 0.10122
Year fixed effects $\checkmark$ $\checkmark$
Country fixed effects $\checkmark$ $\checkmark$
@@ 7 Net output
Table 5: $\theta^{*}_{s}$ and $\theta^{*}_{c}$ in 86 countries (screen = 0.01). Number of years clearing screen shown in ()
Country Period $\overline{\theta^{*}_{s,i}}$ $\overline{\theta^{*}_{c,i}}$ Armenia 1996 - 2018 (21) -0.64 8.04 Aruba 1996 - 2001 (6) 1.21 -4.56 Australia 1980 - 2018 (33) 0.24 15.71 Austria 1996 - 2019 (22) 0.82 27.43 Azerbaijan 1996 - 2018 (21) 1.20 4.94 Bahrain 2009 - 2013 (5) -6.03 -14.91 Belgium 1996 - 2019 (20) 0.61 40.43 Bolivia 1998 - 2015 (17) 0.35 17.13 Botswana 1996 - 1999 (4) 3.14 9.53 Brazil 1996 - 2018 (20) 0.24 6.77 British Virgin Islands 1996 - 1999 (4) 2.39 -1.95 Bulgaria 1996 - 2016 (17) 0.08 20.13 Burkina Faso 2000 - 2018 (19) 0.21 11.07 Cameroon 1997 - 2003 (7) 0.94 -3.99 Canada 1980 - 2020 (36) 0.39 11.42 Cape Verde 2008 - 2017 (9) 0.55 -0.71 Chile 1997 - 2018 (18) 1.06 8.96 China 1992 - 2016 (25) 0.84 11.32 Colombia 1996 - 2019 (24) 0.71 -0.93 Costa Rica 2013 - 2016 (4) 0.62 14.08 Croatia 1996 - 2019 (21) 0.73 -1.55 Curaçao 2001 - 2016 (15) 1.00 5.15 Cyprus 1996 - 2019 (23) 0.12 4.81 Czechia 1994 - 2019 (18) 0.22 11.08 Côte d’Ivoire 1996 - 2000 (5) 0.36 -5.14 Denmark 1996 - 2020 (24) 0.26 2.83 Dominican Republic 1996 - 2016 (10) 1.43 1.55 Ecuador 2008 - 2018 (11) 0.59 -0.60 Egypt 1997 - 2015 (19) 0.95 -1.80 Estonia 1996 - 2019 (22) 0.34 11.35 Finland 1996 - 2020 (19) 0.29 11.85 France 1980 - 2019 (31) 0.23 6.70 Germany 1980 - 2020 (36) 0.70 21.93 Greece 1996 - 2019 (23) 0.10 0.95 Guatemala 2002 - 2019 (18) -0.70 8.22 Guinea 2004 - 2010 (5) 0.75 14.84 Honduras 2001 - 2015 (14) -0.40 5.61 Hong Kong 1997 - 2020 (23) 0.79 13.54 Hungary 1996 - 2019 (21) 0.11 11.65 Iceland 2001 - 2014 (14) 0.20 7.65 India 1999 - 2017 (19) 0.57 10.09 Iran 1997 - 2018 (21) 5.03 4.03 Ireland 1996 - 2019 (23) 0.44 8.00 Country Period $\overline{\theta^{*}_{s,i}}$ $\overline{\theta^{*}_{c,i}}$ Israel 1996 - 2019 (24) 0.44 7.08 Italy 1980 - 2020 (35) 0.11 -3.71 Japan 1980 - 2017 (27) 0.18 7.45 Kazakhstan 1997 - 2019 (22) 0.50 6.01 Kuwait 2003 - 2017 (14) 1.14 8.78 Kyrgyzstan 1996 - 2019 (23) 0.73 -1.01 Latvia 1996 - 2019 (23) -0.21 15.09 Lithuania 1996 - 2019 (22) 0.11 12.21 Luxembourg 1996 - 2017 (20) 0.51 13.46 Malaysia 2007 - 2015 (9) 2.27 1.18 Malta 1996 - 2019 (23) 0.35 25.23 Mexico 1996 - 2019 (20) 0.27 12.74 Moldova 1996 - 2019 (21) -0.62 2.00 Mongolia 2006 - 2019 (14) -0.28 -9.53 Morocco 1999 - 2019 (21) 1.69 7.54 Netherlands 1996 - 2019 (23) 0.88 23.79 New Zealand 1996 - 2019 (21) 0.66 14.13 Nicaragua 2007 - 2018 (12) 0.08 7.53 Niger 1996 - 2019 (24) 0.90 2.10 Norway 1981 - 2020 (37) 0.87 13.94 Peru 2008 - 2019 (12) 0.83 15.55 Philippines 1996 - 2019 (24) 1.53 -2.02 Poland 1996 - 2019 (23) 0.61 10.15 Portugal 1996 - 2020 (22) 0.01 16.81 Qatar 2002 - 2017 (14) 1.85 2.18 Romania 1996 - 2019 (22) 0.11 12.56 Russia 1996 - 2017 (12) 0.02 -2.70 Saudi Arabia 2003 - 2009 (7) 2.69 13.89 Serbia 1998 - 2019 (17) -0.45 -0.28 Slovakia 1996 - 2019 (22) 0.20 17.95 Slovenia 1996 - 2019 (20) 0.32 13.84 South Africa 1996 - 2020 (22) 0.31 18.12 South Korea 1996 - 2018 (22) 0.66 19.05 Spain 1995 - 2019 (22) 0.38 17.37 Sweden 1980 - 2020 (37) 0.25 5.03 Switzerland 1993 - 2019 (23) 0.79 21.89 Tunisia 1996 - 2011 (16) 0.29 -1.58 Turkey 2010 - 2017 (8) 1.23 2.47 USA 1980 - 2018 (34) 0.34 17.53 Ukraine 1996 - 2019 (21) 0.20 2.38 United Kingdom 1982 - 2017 (32) 0.17 15.72 Vanuatu 2002 - 2007 (6) 0.33 9.95 Venezuela 1998 - 2019 (21) 0.65 0.41
Note: Thrift theory predicts $\overline{\theta^{*}_{s,i}}\cong 1$ and $\overline{\theta^{*}_{c,i}}\cong 1$. Free growth theory makes no prediction for these data.
@@ 8 Data sources
All our data are drawn from Distributional National Accounts (DINA) from the free online database World Inequality Database (WID). This source collates data from national accounts and tax data of 105 countries in constant currency units, and adjusts them where needed to conform to current standards of the System of National Accounts (SNA) published by the United Nations. We show results for the 86 of those countries which report all three of the factors, namely net saving, consumption and market-value capital, needed for deriving the thrift and free growth indexes. WID’s source for these data is national accounts.
Consumption $C$ in our text and equations is reproduced from the sum of Government Final Consumption Expenditure (mcongo) [^5] WID code and Private Expenditures of Households and NPISH (mconhn). This sums government expenditure GCE and personal consumption expenditure PCE. Net saving $S_{net}$ and market-value $K$ are taken from net national saving (msavin) and market-value Capital Wealth (mnweal) respectively. GDP, which we use only for weighting purposes in Figs. 1 and 2, is reproduced from GDP (mgdpro).
@@ 9 Accessing our results and methods
Tables and other displays of our findings for each country, and showing our methods of calculation, can be accessed at the web appendix (https://web-appendix.shinyapps.io/Sources_of_Capital_Growth).
@@ 10 Displays
Eqs. (10) and (19) show $\theta_{s}+\varphi_{s}=1$ and $\theta_{c}+\varphi_{c}=1$. Some displays here and in the web appendix save space by showing $\varphi_{s}$ and $\varphi_{c}$ only, leaving $\theta_{s}$ and $\theta_{c}$ implicit as their complements to unity. The web appendix includes displays of $\varphi_{s}$ and $\varphi_{c}$ for each year in each country over the report period. These tend to show upward and downward spikes in values of $\varphi_{s}$ and $\varphi_{c}$ in some years. Those spikes tend to be associated with small absolute values of denominators, in these cases $\Delta g(K)$, in those countries and years. Small denominators magnify errors in measurements of numerators. Worse, when denominators are small, small mismeasurements in them might reverse the term in sign.
To maximize reliability of test results, we apply a range of screens to omit years where absolute denominators fall below a given threshold. Some displays show data for all years, regardless of denominator size. Others screen out all years where absolute denominators are less then .01, then .025, then continuing upward in increments of .025 to a maximum screen of .15. $\varphi_{s}$ and $\varphi_{c}$ are plotted for each country unscreened and at each of the seven successive levels of screening. All displays applied a screen of .01. The denominator whose absolute value is screened is capital acceleration $\Delta g(K)$ or capital growth rate $g(K)$ in all displays.
@@ 12 Discussion and conclusions
These findings challenge the teachings that capital growth is effected by net saving enabled by consumption restraint, and that producer cost, including imputed interest as the opportunity cost of capital, converges to market realization. Evidence showing $\overline{\varphi_{s,i}}\cong 1$ and $\overline{\varphi_{c,i}}\cong 1$ suggests that all capital growth is free, and consequently that market realization, in the presence of innovation, exceeds producer cost by the entirety of capital growth. Meanwhile the same evidence, which indicates that net saving adds no capital value, suggests a review of the teachings that consumption plus net saving gives net income, and that consumption plus net investment gives net output. Appendix B will also question the latter teachings.
@@ Appendix B.  Net output with human capital
It is possible in principle to model a value-added chain which includes human capital, and to compare findings with those shown in Eq. (B.1). Let human capital $H$, in that new model, stand as the last link in the value-added chain. Adapting the classic illustration of the value added principle, say that farms produce wheat, mills convert the wheat to flour, bakeries convert the flour into bread, and humans convert some of the bread, called invested consumption, into human capital. The net outflow from this extended value-added chain is not all of consumption, but only pure consumption, meaning the part remaining after the part invested in human capital (invested consumption) is subtracted [^7] The concepts of invested consumption, pure consumption, self-invested work and human depreciation were introduced in Schultz (1961).. By this reasoning, the principle that net output is expressed in capital growth plus net outflow gives
@@
if Schultz and Ben-Porath and the reasoning here are right.
@@ Appendix C.  Mill’s statement of the free growth idea
This passage may be the first clear statement of what we call free growth theory. His use of the words "accelerate" and "concurrently" suggest that his path of reasoning was something like ours, and that he meant what we call free growth rather than the alternating phases of higher and lower saving rates pictured in thrift theory.